OpenStudy (anonymous):

Find the product 9.9.9.9

OpenStudy (anonymous):

? 9*9*9*9 ? 9*9*9*9= 9^4 9*9=81 81*81=?

OpenStudy (anonymous):

6561

OpenStudy (anonymous):

\[9\times9\times9\times9=9^4\] 9x9=81 81x9=729 729x9=6561

OpenStudy (anonymous):

6561